# Symbiotic Backscatter Communication: A Design Perspective on the Modulation Scheme of Backscatter Devices

### v2

@@ I Introduction
In this paper, we focus on the SRBC, where the BD [^2] A reconfigurable intelligent surface (RIS) is typically designed to enhance wireless transmission performance by optimizing channel conditions, whereas a backscatter device (BD) is aimed at delivering information with ultra-low device complexity and ultra-low power consumption. It is worth noting that recent research has explored the possibility of enabling RIS to modulate information [27], suggesting that an RIS could, in principle, replace the BD in our considered system model. However, the hardware complexity and manufacturing cost of RIS units are substantially higher than those of BDs. Therefore, RIS may not be a suitable solution for large-scale deployments of ultra low-power and low-cost IoT devices. This practical constraint motivates our focus on the use of BDs in this work. adopts one of the following two popular low-complexity modulation schemes, i.e., $M$-ary amplitude-shift keying (MASK) and $M$-ary phase-shift keying (MPSK) [^3] MASK and MPSK are selected for their dual advantages of being recommended by 3GPP for IoT and enabling simple, low-power passive implementations suitable for BDs [3]., to convey information. It should be noted that under MASK/MPSK, the BD’s signal does not follow a CSCG distribution. Consequently, the following questions arise: First, can BD still guarantee an increase in the PT’s rate when using MASK or MPSK? If not, how should the phase of the BD’s symbols be designed to ensure an improvement in the PT’s rate? Second, the symbols generated by the modulation scheme do not follow CSCG distribution, and applying Shannon formula would lead to an overestimation of the achievable rate. In this case, what is the appropriate method for calculating the BD’s rate? Although the recent work [28] employs low-complexity modulation schemes, it does not address the above questions. Our work aims to fill this gap by providing a practical and theoretically grounded design framework for SRBC under realistic modulation constraints.
@@
  • Simulation results confirm that the optimal phase of MASK/MPSK can improve the PT’s rate, and demonstrate that the proposed discrete-phase circuit design achieves near-optimal performance. Besides, the results also reveal the condition under which ASK outperforms PSK in terms of the BD’s rate.
@@ II System Model
Fig. 1 depicts a SBC system, where the receiver (R) simultaneously extracts the PT’s information and the BD’s information that is modulated onto the PT’s signal. Let $h_{1}$, $h_{2}$, and $h_{3}$ denote the channel coefficients of the PT-R link, the PT-BD link, and the BD-R link, respectively, where $|h_{i}|$ and $\theta_{i}$ $\left(i=1,2,3\right)$ represent the amplitude and the phase of $h_{i}$, respectively. All the channel coefficients are assumed to remain fixed within a transmission block but may change across different blocks. Let $s\left(n\right)$ denote the signal transmitted by the PT, which has a mean of zero and a variance of $P$, with a symbol period of $T_{s}$. The period of a BD symbol is denoted by $T_{c}$. Since the period of a BD symbol is much longer than that of the PT, we assume $T_{c}=LT_{s}$, where $L\gg 1$ is a positive integer [9, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
@@ III-A MASK Modulation
In what follows, we derive the PT’s rate by assuming $M\to\infty$. Under this assumption, $\alpha_{m}$ follows the continuous uniform distribution from 0 to 1, and the PT’s rate can be calculated as (III-A), as shown at the top of the next page, where ${C_{1}}=\frac{P}{{{\sigma^{2}}}}{\left|{{h_{1}}}\right|^{2}}+1$, $C_{2}=2\frac{P}{{{\sigma^{2}}}}\left|{{h_{1}}}\right|\left|{{h_{2}}}\right|\left|{{h_{3}}}\right|\cos\left({\theta_{0}+\varphi_{0}^{\rm{A}}}\right)$ and $C_{3}=\frac{P}{{{\sigma^{2}}}}{\left|{{h_{2}}}\right|^{2}}{\left|{{h_{3}}}\right|^{2}}$, $\theta_{0}=\theta_{2}+\theta_{3}-\theta_{1}$. In (III-A), the second equality holds from the integration by parts, and the last equality is derived by using $\int\frac{{\left({{N_{1}}x+{N_{2}}}\right)dx}}{{E+2Fx+G{x^{2}}}}=\frac{M}{{2C}}\ln\left|{E+2Fx+G{x^{2}}}\right|+\frac{{{N_{2}}G-{N_{1}}F}}{{G\sqrt{EG-{F^{2}}}}}\arctan\frac{{Gx+F}}{{\sqrt{EG-{F^{2}}}}}$ when $EG>{F^{2}}$ [33, eq. (2.103.5)], and ${C_{1}}{C_{3}}-{\left({\frac{{{C_{2}}}}{2}}\right)^{2}}=\frac{{{P^{2}}}}{{{\sigma^{4}}}}{\left|{{h_{1}}}\right|^{2}}{\left|{{h_{2}}}\right|^{2}}{\left|{{h_{3}}}\right|^{2}}\left({1-\cos\left(\theta_{0}+\varphi_{0}^{\rm{A}}\right)}\right)+\frac{P}{{{\sigma^{2}}}}{\left|{{h_{2}}}\right|^{2}}{\left|{{h_{3}}}\right|^{2}}>0$.
@@
$={\log_{2}}\left({{C_{1}}+{C_{2}}+{C_{3}}}\right)-\frac{2}{{\ln 2}}+\frac{{{C_{2}}}}{{2{C_{3}}\ln 2}}\ln\frac{{\left|{{C_{1}}}\right|}}{{\left|{{C_{1}}+{C_{2}}+{C_{3}}}\right|}}+\frac{{2{C_{1}}-\frac{{{{\left({{C_{2}}}\right)}^{2}}}}{2}}}{{{C_{3}}\sqrt{{C_{3}}{C_{1}}-{{\left({\frac{{{C_{2}}}}{2}}\right)}^{2}}}\ln 2}}$
$\times\left({\arctan\frac{{{C_{2}}}}{{\sqrt{{C_{3}}{C_{1}}-{{\left({\frac{{{C_{2}}}}{2}}\right)}^{2}}}}}-\arctan\frac{{{C_{3}}+{C_{2}}}}{{\sqrt{{C_{3}}{C_{1}}-{{\left({\frac{{{C_{2}}}}{2}}\right)}^{2}}}}}}\right)$ (11)
@@ VI-1 Example 1
In this example, we assume that there exist $\varphi_{0}^{\rm P}$} and $\varphi_{0}$ satisfying $\theta_{2}+\theta_{3}-\theta_{1}+{\varphi_{0}^{\rm P}}=0$ and $\left|{{h_{1}}}\right|={\alpha_{0}^{\rm P}}\left|{{h_{2}}}\right|\left|{{h_{3}}}\right|$, and prove that the rate gain of the PT under the above assumption is lower than zero. Based on this assumption, according to (13), $R^{\rm{P}}_{\rm s}$ can be rewritten as
@@
Then, we can write $\Delta{R_{s}}$ as
@@
$\Delta{R_{\rm s}^{\rm{P}}}={\log_{2}}\left({\frac{{\prod\limits_{m=1,m\neq{m_{0}}}^{M}{{{\left({1+2b\left({1+\cos\left({\frac{{2\pi}}{M}\left({m-1}\right)}\right)}\right)}\right)}^{\frac{1}{M}}}}}}{{1+b}}}\right).$ (27)
@@ VI-2 Example 2
In this example, we assume that there exist $\varphi_{0}^{\rm P}$} and $\alpha_{0}$ meeting $\theta_{2}+\theta_{3}-\theta_{1}+{\varphi_{0}^{\rm P}}=\frac{\pi}{2}$ and $\left|{{h_{1}}}\right|={\alpha_{0}^{\rm{P}}}\left|{{h_{2}}}\right|\left|{{h_{3}}}\right|$, and prove that the rate gain of the PT under the above assumption is larger than zero. $\Delta{R_{\rm s}^{\rm{P}}}$ can be rewritten as

### v3

@@ I Introduction
In this paper, we focus on the SBC, where the BD [^2] A reconfigurable intelligent surface (RIS) is typically designed to enhance wireless transmission performance by optimizing channel conditions, whereas a backscatter device (BD) is aimed at delivering information with ultra-low device complexity and ultra-low power consumption. It is worth noting that recent research has explored the possibility of enabling RIS to modulate information [27], suggesting that an RIS could, in principle, replace the BD in our considered system model. However, the hardware complexity and manufacturing cost of RIS units are substantially higher than those of BDs. Therefore, RIS may not be a suitable solution for large-scale deployments of ultra low-power and low-cost IoT devices. This practical constraint motivates our focus on the use of BDs in this work. adopts one of the following two popular low-complexity modulation schemes, i.e., $M$-ary amplitude-shift keying (MASK) and $M$-ary phase-shift keying (MPSK) [^3] MASK and MPSK are selected for their dual advantages of being recommended by 3GPP for IoT and enabling simple, low-power passive implementations suitable for BDs [3]., to convey information. It should be noted that under MASK/MPSK, the BD’s signal does not follow a CSCG distribution. Consequently, the following questions arise: First, can BD still guarantee an increase in the PT’s rate when using MASK or MPSK? If not, how should the phase of the BD’s symbols be designed to ensure an improvement in the PT’s rate? Second, the symbols generated by the modulation scheme do not follow CSCG distribution, and applying Shannon formula would lead to an overestimation of the achievable rate. In this case, what is the appropriate method for calculating the BD’s rate? Although the recent work [28] employs low-complexity modulation schemes, it does not address the above questions. Our work aims to fill this gap by providing a practical and theoretically grounded design framework for SRBC under realistic modulation constraints.
@@
  • Simulation results confirm that the optimal phase of MASK/MPSK can improve the PT’s rate, and demonstrate that the proposed discrete-phase circuit design achieves near-optimal performance. Besides, the results also reveal the condition under which ASK outperforms PSK in terms of the PT’s rate.
@@ II System Model
Fig. 1 depicts a SBC system, where the receiver (R) simultaneously extracts the PT’s information and the BD’s information that is modulated onto the PT’s signal. Let $h_{1}$, $h_{2}$, and $h_{3}$ denote the channel coefficients of the PT-R link, the PT-BD link, and the BD-R link, respectively, where $|h_{i}|$ and $\theta_{i}$ $\left(i=1,2,3\right)$ represent the amplitude and the phase of $h_{i}$, respectively. All the channel coefficients are assumed to remain fixed within a transmission block but may change across different blocks. Let $s\left(n\right)$ denote the signal transmitted by the PT, which has a mean of zero and a variance of one, with a symbol period of $T_{s}$. The period of a BD symbol is denoted by $T_{c}$. Since the period of a BD symbol is much longer than that of the PT, we assume $T_{c}=LT_{s}$, where $L\gg 1$ is a positive integer [9, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
@@ III-A MASK Modulation
In what follows, we derive the PT’s rate by assuming $M\to\infty$. Under this assumption, $\alpha_{m}$ follows the continuous uniform distribution from 0 to 1, and the PT’s rate can be calculated as (III-A), as shown at the top of the next page, where ${C_{1}}=\frac{P}{{{\sigma^{2}}}}{\left|{{h_{1}}}\right|^{2}}+1$, $C_{2}=2\frac{P}{{{\sigma^{2}}}}\left|{{h_{1}}}\right|\left|{{h_{2}}}\right|\left|{{h_{3}}}\right|\cos\left({\theta_{0}+\varphi_{0}^{\rm{A}}}\right)$ and $C_{3}=\frac{P}{{{\sigma^{2}}}}{\left|{{h_{2}}}\right|^{2}}{\left|{{h_{3}}}\right|^{2}}$, $\theta_{0}=\theta_{2}+\theta_{3}-\theta_{1}$. In (III-A), the second equality holds from the integration by parts, and the last equality is derived by using $\int\frac{{\left({{N_{1}}x+{N_{2}}}\right)dx}}{{E+2Fx+G{x^{2}}}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{N_{1}}{{2G}}}\ln\left|{E+2Fx+G{x^{2}}}\right|+\frac{{{N_{2}}G-{N_{1}}F}}{{G\sqrt{EG-{F^{2}}}}}\arctan\frac{{Gx+F}}{{\sqrt{EG-{F^{2}}}}}$ when $EG>{F^{2}}$ [33, eq. (2.103.5)], and ${C_{1}}{C_{3}}-{\left({\frac{{{C_{2}}}}{2}}\right)^{2}}=\frac{{{P^{2}}}}{{{\sigma^{4}}}}{\left|{{h_{1}}}\right|^{2}}{\left|{{h_{2}}}\right|^{2}}{\left|{{h_{3}}}\right|^{2}}\left({1-\cos^{2}\left(\theta_{0}+\varphi_{0}^{\rm{A}}\right)}\right)+\frac{P}{{{\sigma^{2}}}}{\left|{{h_{2}}}\right|^{2}}{\left|{{h_{3}}}\right|^{2}}>0$.
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}={\log_{2}}\left({{C_{1}}+{C_{2}}+{C_{3}}}\right)-\frac{2}{{\ln 2}}-\frac{{{C_{2}}}}{{2{C_{3}}\ln 2}}\ln\frac{{\left|{{C_{1}}}\right|}}{{\left|{{C_{1}}+{C_{2}}+{C_{3}}}\right|}}-\frac{{2{C_{3}}{C_{1}}-\frac{{{{\left({{C_{2}}}\right)}^{2}}}}{2}}}{{{C_{3}}\sqrt{{C_{3}}{C_{1}}-{{\left({\frac{{{C_{2}}}}{2}}\right)}^{2}}}\ln 2}}}$
$\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left({\arctan\frac{{\frac{{{C_{2}}}}{2}}}{{\sqrt{{C_{3}}{C_{1}}-{{\left({\frac{{{C_{2}}}}{2}}\right)}^{2}}}}}-\arctan\frac{{{C_{3}}+\frac{{{C_{2}}}}{2}}}{{\sqrt{{C_{3}}{C_{1}}-{{\left({\frac{{{C_{2}}}}{2}}\right)}^{2}}}}}}\right)}$ (11)
@@ VI-1 Example 1
In this example, we assume that there exist $\varphi_{0}^{\rm P}$} and $\alpha_{0}^{\rm P}$} satisfying $\theta_{2}+\theta_{3}-\theta_{1}+{\varphi_{0}^{\rm P}}=0$ and $\left|{{h_{1}}}\right|={\alpha_{0}^{\rm P}}\left|{{h_{2}}}\right|\left|{{h_{3}}}\right|$, and prove that the rate gain of the PT under the above assumption is lower than zero. Based on this assumption, according to (13), $R^{\rm{P}}_{\rm s}$ can be rewritten as
@@
Then, we can write $\Delta{R_{s}}^{\rm{P}}$ as
@@
$\Delta{R_{\rm s}^{\rm{P}}}={\log_{2}}\left({\frac{{\prod\limits_{m=1,m\neq{m_{0}}}^{M}{{{\left({1+2b\left({1+\cos\left({\frac{{2\pi}}{M}\left({m-1}\right)}\right)}\right)}\right)}^{\frac{1}{M}}}}}}{{1+b}}}\right).$ (C.4)
@@ VI-2 Example 2
In this example, we assume that there exist $\varphi_{0}^{\rm P}$} and $\alpha_{0}^{\rm P}$ meeting $\theta_{2}+\theta_{3}-\theta_{1}+{\varphi_{0}^{\rm P}}=\frac{\pi}{2}$ and $\left|{{h_{1}}}\right|={\alpha_{0}^{\rm{P}}}\left|{{h_{2}}}\right|\left|{{h_{3}}}\right|$, and prove that the rate gain of the PT under the above assumption is larger than zero. $\Delta{R_{\rm s}^{\rm{P}}}$ can be rewritten as